# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v3

@@ Abstract
Text embedding has become a foundational technology in natural language processing (NLP) during the deep learning era, driving advancements across a wide array of downstream tasks. While many natural language understanding challenges can now be modeled using generative paradigms and leverage the robust generative and comprehension capabilities of large language models (LLMs), numerous practical applications—such as semantic matching, clustering, and information retrieval—continue to rely on text embeddings for their efficiency and effectiveness. Therefore, integrating LLMs with text embeddings has become a major research focus in recent years. In this survey, we categorize the interplay between LLMs and text embeddings into three overarching themes: (1) LLM-augmented text embedding, enhancing traditional embedding methods with LLMs; (2) LLMs as text embedders, adapting their innate capabilities for high-quality embedding; and (3) Text embedding understanding with LLMs, leveraging LLMs to analyze and interpret embeddings. By organizing recent works based on interaction patterns rather than specific downstream applications, we offer a novel and systematic overview of contributions from various research and application domains in the era of LLMs. Furthermore, we highlight the unresolved challenges that persisted in the pre-LLM era with pre-trained language models (PLMs) and explore the emerging obstacles brought forth by LLMs. Building on this analysis, we outline prospective directions for the evolution of text embedding, addressing both theoretical and practical opportunities in the rapidly advancing landscape of NLP.
@@ 2.1.4 Era of Pre-trained Language Models
where $s:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the distance function, while $x^{+}$ and $\{x^{-}_{j}\}_{j=1}^{N}$ is the positive example and negative examples for the text $x$, respectively. Improvements based on the contrastive learning have focused on (a) the form of the loss function [81, 82, 83, 84] and (b) better methods for constructing positive and negative samples [79, 85, 86, 80]. For example, ANCE [87] explores a two-stage training method. First, it utilizes non-embedding model methods, such as BM25, to extract a batch of negative samples for model training. Subsequently, it employs the trained embedding model to re-mine higher-quality negative samples for further model training. This method has become a standard practice for training embedding models. Conan-Embedding [88] employs a dynamic hard negative mining method to maximize the utilization of high-quality negative examples. However, real-world datasets often suffer from incomplete positive sample labeling due to cost constraints; for most datasets, only around one positive sample is labeled per query. This necessitates careful consideration of the influence of false negatives during negative sample mining. Many researchers suggest that sampling from the middle of the embedding retrieval results or filtering based on similarity scores can effectively enhance model performance [89, 90]. (3) Pre-training paradigm. Contrastive learning is useful on large-scale corpora but takes big performance hits on low data situations [74]. This observation shows that the pre-training tasks of PLMs, i.e., Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), etc., do not allow the PLMs to adapt to semantic aggregation quickly and output high-quality text embeddings. Therefore, some work has been devoted to designing better pre-training tasks to improve the performance of contrastive learning in low-data scenarios. Early explorations [91, 92] used only additional incremental pre-training tasks, such as the Inverse Cloze Task (ICT), to improve contrast learning performance. Subsequent improvements involve the modification of model architectures. For example, Condenser [93] restores the masked token using the [CLS]’s hidden states from the last layer and other tokens’ hidden states from a shallow layer. Therefore, the newly generated information in the deep layers has to aggregate in the [CLS] to reconstruct the masked information. SEED-Encoder [94] and RetroMAE [95] feed the final hidden state of [CLS] into an auxiliary decoder to restore the masked information, encouraging the aggregation of information at [CLS].
@@ 2.4.1 Long Context Compression
[…]
@@ 2.4.2 Embedding Inversion
    Cosine $\text{Cosine Similarity}=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$ (17)
@@
[…]
@@ 4 LLMs as Text Embedder
4.1 Backbone Selection
Model
The LLM-based embedders appearing in Table 4 use numerous open-source LLMs as backbones. Among the encoder-decoder backbones, T5 (GTR) has a dominant position, which is not surprising, as T5 and its variants have been leading innovations in encoder-decoder PLMs. Among decoder-only backbones, the most popular is Mistral (20 / 40), followed by LLaMA (17 / 40) and Qwen (6 / 40). The popularity of the Mistral is strongly connected to its excellent performance: a recent study [168] find that Mistral’s embedding performance will not drop significantly compared to the other LLMs when converting causal attention to bi-directional attention. This characteristic may derive from its pre-training or instruction fine-tuning phase, but no information has been revealed.
Size
4.2 Architecture Improvement
Denoted $d$ is the dimension of hidden states, $L$ is the number of transformer layers, and $c$ is the vocabulary size. Without loss of generality, we divide the LLM $F$ into multiple Transformer layers and a decoding layer, which is denoted as
$F=g\circ[f^{(L)}\circ\cdots\circ f^{(1)}]\circ f^{(0)}=g\circ f$
where $f^{(0)}:\mathcal{X}\to\mathbb{R}^{d}$ is the input layer, $f^{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the $l$-th transformer layer and $g:\mathbb{R}^{d}\to\mathbb{R}^{c}$ is the decoding layer. For convenience, we use $f$ to represent $[f^{(L)}\circ\cdots\circ f^{(1)}]\circ f^{(0)}$.
4.2.1 Pooling Strategy
Similar to traditional language models, there is a decoding layer within the LLM to complete the mapping from the hidden state space to the token vocabulary space. To the best of our knowledge, almost all LLMs follow the original design of GPT [171], using a simple unbiased linear layer as the decoding layer. When LLMs are used as textual embedders, embeddings are obtained by a specific pooling strategy at hidden states, while mapping to token space is no longer necessary. Therefore, most methods directly discard the decoding layer and opt for a pooling strategy on the hidden state of the last transformer layer’s output to obtain a single embedding for each text.
$[{\bm{h}}^{(L)}_{1},\cdots{\bm{h}}^{(L)}_{n}]=\mathcal{H}=f(x),\quad x\in\mathcal{X},$
[…]
where $n$ is the token number of the input $x$ depended on the tokenizer of the LLMs, and $\{\alpha_{i}\}_{i=1}^{n}$ is the weight satisfying $\sum_{i}^{n}\alpha_{i}=1$.
Four pooling methods fit this scenario: (Weighted) Mean Pooling, First Pooling, Last Pooling and Partial Pooling. However, some methods are not simply weighting the hidden states of the last layer. For example, some methods use learnable modules to interact with the hidden states of the last layer (Post-Interaction Pooling), while others try to augment the final embedding with other information (e.g., attentional weights, hidden states, etc.) that utilizes the output of the middle layer (Multi-Layer Pooling). We cover them one by one below.
(Weighted) Mean Pooling
Mean pooling ($\alpha_{i}=\frac{1}{n}$ for each $i$) averages the embedding of each position and shows better performance in BERT and T5 than first pooling [72, 172]. Meanwhile, in decoder-based LLMs, it is not reasonable to assign the same weight to each position because the latter position will be able to see more semantic information due to causal attention [173]. Therefore, SGPT [173] suggests the use of weighted mean pooling ($\alpha_{i}=\frac{i}{\sum_{i=1}^{n}i}$) with the intuition that the latter position should be given larger weight. L3Prune [174] follows SGPT’s pooling strategy to fine-tune the pruned LLM-based embedders.
Table 4: The Overview of LLMs as Text Encoders. The paradigm part is shortened as follows: supervised contrastive learning (SCL), weak supervised contrastive learning (WCL), unsupervised contrastive learning (UCL), training-free (TF), next-token prediction (NTP), Generation-augmented Representation Learning (GRL), FLOPS regularization (FLOPs), masked next-token prediction (MNTP) and In-Context Learning (ICL).
@@ 4.2.1 Pooling Strategy
Sentence-T5 [172] T5 Frist / Mean Bi $\surd$ $\times$ - WCL→SCL ZS / FT STS / Clf.
@@
BeLLM [178] LLaMA Last Bi $\times$ LoRA Prompt SCL ZS / FT STS / Clf.
GTR [179] T5 Mean Bi $\surd$ $\times$ - WCL→SCL ZS / FT IR
SGPT-IR [173] GPT-Neo Weighted Mean Causal $\times$ BitFit - SCL ZS / FT IR
Llama2Vec [180] LLaMA Special Last Causal $\times$ LoRA - EBAE+EBAR→SCL FT IR
@@
BMRetriever [182] Pythia, Gemma, BioMistral Special Last Causal $\times$ LoRA Instruction WCL→SCL ZS / FT IR
@@
LMORT [185] GPT2, GPT-J Multi-Layer Causal $\surd$ $\times$ - SCL FT IR
Mistral-SPLADE [186] Mistral Post Causal $\times$ QLoRA Prompt SCL+FLOPS ZS IR
NV-Retriever [153] Mistral Mean Bi $\times$ LoRA Instruction SCL→SCL ZS IR
@@
RARe [187] LLaMA, Mistral Mean / Last Bi / Causal $\times$ LoRA Instruction+Example SCL ICL IR
DEBATER [188] MiniCPM Special Seq Last Causal $\times$ LoRA Instruction SCL+SD ZS IR
O1 Embedder [189] Mistral, LLaMA, Qwen Special Last Causal $\times$ LoRA Instruction SCL+NTP ZS IR
@@
E5-Mistral [157] Mistral Last Causal $\times$ LoRA Instruction SCL ZS Uni.
SFR-Mistral [194] Mistral Last Causal $\times$ LoRA Instruction SCL ZS Uni.
LLM2Vec [168] LLaMA, Mistral Mean Bi $\times$ LoRA Instruction MNTP→UCL/SCL ZS Uni.
@@
UDEVER [196] BLOOM Special Last Causal $\times$ BitFit - SCL ZS Uni.
InstructOR [67] GTR Mean Bi $\surd$ $\times$ Instruction SCL ZS Uni.
InBedder [197] OPT, LLaMA Last Causal $\times$ $\times$ Instruction+Prompt NTP ZS Uni.
GritLM [169] Mistral Mean Customized $\surd$ / $\times$ $\times$ Instruction SCL+NTP ZS / FS Uni.
GTE-Qwen2 [198] Qwen Last Bi $\times$ $\times$ Instruction WCL→SCL ZS Uni.
NV-Embed [199] Mistral Post Bi $\times$ LoRA Instruction WCL→SCL ZS Uni.
MLTP [200] Mistral Multi-Layer Bi / Causal $\times$ $\times$ Instruction SCL ZS Uni.
ULLME [201] LLaMA, Mistral, Phi Mean Bi / Causal $\times$ LoRA - GRL ZS Uni.
@@
BGE-ICL [202] Mistral, Gemma Special Last Bi / Causal $\times$ LoRA Instruction+Example SCL ZS / ICL Uni.
@@
Linq-Embed [203] Mistral Last Causal $\times$ LoRA Instruction SCL ZS Uni.
@@
LENS [204] Mistral Post Bi $\times$ LoRA Instruction SCL ZS Uni.

### v4

@@ Abstract
Text embedding has become a foundational technology in natural language processing (NLP) during the deep learning era, driving advancements across a wide array of downstream tasks. While many natural language understanding challenges can now be modeled using generative paradigms and leverage the robust generative and comprehension capabilities of large language models (LLMs), numerous practical applications, such as semantic matching, clustering, and information retrieval, continue to rely on text embeddings for their efficiency and effectiveness. Therefore, integrating LLMs with text embeddings has become a major research focus in recent years. In this survey, we categorize the interplay between LLMs and text embeddings into three overarching themes: (1) LLM-augmented text embedding, enhancing traditional embedding methods with LLMs; (2) LLMs as text embedders, adapting their innate capabilities for high-quality embedding; and (3) Text embedding understanding with LLMs, leveraging LLMs to analyze and interpret embeddings. By organizing recent works based on interaction patterns rather than specific downstream applications, we offer a novel and systematic overview of contributions from various research and application domains in the era of LLMs. Furthermore, we highlight the unresolved challenges that persisted in the pre-LLM era with pre-trained language models (PLMs) and explore the emerging obstacles brought forth by LLMs. Building on this analysis, we outline prospective directions for the evolution of text embedding, addressing both theoretical and practical opportunities in the rapidly advancing landscape of NLP.
@@ 2.1.4 Era of Pre-trained Language Models
where $s:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the distance function, while $x^{+}$ and $\{x^{-}_{j}\}_{j=1}^{N}$ is the positive example and negative examples for the text $x$, respectively.
Technical Improvements
Improvements based on the contrastive learning have focused on (a) the form of the loss function [81, 82, 83, 84] and (b) better methods for constructing positive and negative samples [79, 85, 86, 80]. For example, ANCE [87] explores a two-stage training method. First, it utilizes non-embedding model methods, such as BM25, to extract a batch of negative samples for model training. Subsequently, it employs the trained embedding model to re-mine higher-quality negative samples for further model training. This method has become a standard practice for training embedding models. Conan-Embedding [88] employs a dynamic hard negative mining method to maximize the utilization of high-quality negative examples. However, real-world datasets often suffer from incomplete positive sample labeling due to cost constraints; for most datasets, only around one positive sample is labeled per query. This necessitates careful consideration of the influence of false negatives during negative sample mining. Many researchers suggest that sampling from the middle of the embedding retrieval results or filtering based on similarity scores can effectively enhance model performance [89, 90]. (3) Pre-training paradigm. Contrastive learning is useful on large-scale corpora but takes big performance hits on low data situations [74]. This observation shows that the pre-training tasks of PLMs, i.e., Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), etc., do not allow the PLMs to adapt to semantic aggregation quickly and output high-quality text embeddings. Therefore, some work has been devoted to designing better pre-training tasks to improve the performance of contrastive learning in low-data scenarios. Early explorations [91, 92] used only additional incremental pre-training tasks, such as the Inverse Cloze Task (ICT), to improve contrast learning performance. Subsequent improvements involve the modification of model architectures. For example, Condenser [93] restores the masked token using the [CLS]’s hidden states from the last layer and other tokens’ hidden states from a shallow layer. Therefore, the newly generated information in the deep layers has to aggregate in the [CLS] to reconstruct the masked information. SEED-Encoder [94] and RetroMAE [95] feed the final hidden state of [CLS] into an auxiliary decoder to restore the masked information, encouraging the information aggregation.
@@ 2.4.1 Long Context Compression
    $\text{ROUGE-L}_{\text{R}}=\frac{\operatorname{LCS}(a_{i},\hat{a}_{i})}{L_{a_{i}}}$ (13)
    $\text{ROUGE-L}_{\text{P}}=\frac{\operatorname{LCS}(a_{i},\hat{a}_{i})}{L_{\hat{a}_{i}}}$ (14)
    $\text{ROUGE-L}=\frac{(1+\beta^{2})\cdot\text{ROUGE-L}_{\text{P}}\cdot\text{ROUGE-L}_{\text{R}}}{\beta^{2}\cdot\text{ROUGE-L}_{\text{P}}+\text{ROUGE-L}_{\text{R}}}$ (15)
@@ 2.4.2 Embedding Inversion
    $\text{Cosine Similarity}=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$ (17)
@@
    $\text{Token F1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (18)
@@ 4 LLMs as Text Embedder
The overview of LLM-based embedders. The Paradigm column is shortened as follows: supervised contrastive learning (SCL), weak supervised contrastive learning (WCL), unsupervised contrastive learning (UCL), training-free (TF), supervised next token prediction (SNTP), unsupervised next token prediction (UNTP), FLOPS regularization (FLOPS), masked next token prediction (MNTP), mask language modeling (MLM), Kullback-Leibler divergence (KL), metric learning (ML), auto-encoding (AE), iterative contrastive refinement (ICR) and reinforcement learning (RL).
@@
[…]
Method Model Architecture Training Evaluation
[…]
@@
BeLLM [173] LLaMA Last Bi-Dir $\times$ LoRA Prompt SCL ZS / FT STS / Clf.
AutoRegEmbed [174] LLaMA, Mistral Special Seq Mean Causal $\times$ $\times$ Instruction SNTP→SCL ZS STS
GTR [175] T5 Mean Bi-Dir $\surd$ $\times$ - WCL→SCL ZS / FT IR
SGPT-IR [176] GPT-Neo Weighted Mean Causal $\times$ Bi-DirtFit - SCL ZS / FT IR
Llama2Vec [177] LLaMA Special Last Causal $\times$ LoRA - AE+UNTP→SCL FT IR
@@
BMRetriever [179] Pythia, Gemma, Bi-DiroMistral Special Last Causal $\times$ LoRA Instruction WCL→SCL ZS / FT IR
@@
LMORT [182] GPT2, GPT-J Multi-Layer Causal $\times$ $\times$ - SCL FT IR
Mistral-SPLADE [183] Mistral Post Causal $\surd$ QLoRA Prompt SCL+FLOPS ZS IR
NV-Retriever [153] Mistral Mean Bi-Dir $\times$ LoRA Instruction SCL→SCL ZS IR
@@
RARe [184] LLaMA, Mistral Mean / Last Bi-Dir / Causal $\times$ LoRA Instruction+Example SCL ICL IR
DEBATER [185] MiniCPM Special Seq Last Causal $\times$ LoRA Instruction SCL+KL ZS IR
O1 Embedder [186] Mistral, LLaMA, Qwen Gen. Seq Last Causal $\times$ LoRA Instruction SCL+SNTP ZS IR
Search-R3 [187] Qwen Gen. Seq Last Causal $\times$ LoRA Instruction SNTP+KL+SCL+ML→RL ZS IR
ReasonEmbed [188] LLaMA, Qwen Special Last Causal $\times$ LoRA Instruction SCL ZS IR
@@
LLM2Vec [193] LLaMA, Mistral Mean Bi-Dir $\times$ LoRA Instruction MNTP→UCL/SCL ZS Uni.
@@
UDEVER [195] BLOOM Special Last Causal $\times$ Bi-DirtFit - SCL ZS Uni.
InstructOR [67] GTR Mean Bi-Dir $\surd$ $\times$ Instruction SCL ZS Uni.
InBedder [196] OPT, LLaMA Last Causal $\times$ $\times$ Instruction+Prompt SNTP ZS Uni.
GritLM [197] Mistral Mean Customized $\surd$ / $\times$ $\times$ Instruction SCL+SNTP ZS / FS Uni.
MLTP [198] Mistral Multi-Layer Bi-Dir / Causal $\times$ $\times$ Instruction SCL ZS Uni.
ULLME [199] LLaMA, Mistral, Phi Mean Bi-Dir / Causal $\times$ LoRA - SCL+RL+KL ZS Uni.
@@
LENS [201] Mistral Post Bi-Dir $\times$ LoRA Instruction SCL ZS Uni.
DIFFEMBED [202] Dream Mean Bi-Dir $\times$ LoRA Instruction SCL ZS Uni.
MGH [203] Mistral Weighted Mean Bi-Dir $\times$ LoRA Instruction SCL ZS Uni.
GRACE [204] LLaMA, Qwen Hyb. Seq Mean Causal $\times$ $\times$ Instruction RL ZS Uni.
@@
Lychee [205] Qwen Special Last Causal $\times$ LoRA Instruction SCL→SCL→MG→SCL ZS Uni.
@@
GIRCSE [206] Mistral, Qwen Gen. Seq Mean Causal $\times$ LoRA Instruction SCL+ICR ZS Uni.
Anchor [207] Mistral, LLaMA, Qwen Special Last Causal $\times$ LoRA Instruction SNTP→SCL ZS Uni.
Text2Token [208] LLaMA, Mistral Mean / Last Bi-Dir / Causal $\times$ LoRA - / Prompt UNTP ZS Uni.
Series works by commercial companies
CoDiEmb (Tecent) [209] MiniCPM Mean Causal $\times$ $\times$ Instruction SCL+Pearson+KL+RL→MG ZS Uni.
Conan-Embedding-v2 (Tecent) [210] 1.4B LLM Mean Causal→Bi-Dir $\times$ $\times$ Instruction UNTP→SNTP→WCL→SCL ZS Uni.
F2LLM (Ant) [211] Qwen Special Last Causal $\times$ $\times$ Instruction SCL ZS Uni.
Linq-Embed-Mistral (Linq AI) [212] Mistral Special Last Causal $\times$ LoRA Instruction SCL ZS Uni.
QZhou-Embedding (KingSoft) [213] Qwen Mean Bi-Dir $\times$ $\times$ Instruction SCL→SCL+ML ZS Uni.
Flag Embedding (BAAI)
BGE-ICL [214] Mistral, Gemma Special Last Bi-Dir / Causal $\times$ LoRA Instruction+Example SCL ZS / ICL Uni.
BGE-Multilingual-Gemma2 Gemma Special Last Causal $\times$ $\times$ Instruction Unknown ZS Uni.
E5 Embedding (Microsoft)
E5-Mistral [157] Mistral Special Last Causal $\times$ LoRA Instruction SCL ZS Uni.
@@
SPEED [215] Mistral Special Last Causal $\times$ LoRA Instruction SCL ZS
Gemini Embedding (Google)
Gecko [156] Unknown Mean Unknown $\times$ $\times$ Instruction WCL→SCL ZS Uni.
Gemini Embedding [216] Gemini Mean Bi-Dir $\surd$ $\times$ Instruction WCL→SCL→MG ZS Uni.
NV-Retriever [153] Mistral Mean Bi-Dir $\times$ LoRA Instruction SCL→SCL ZS IR
NV-Embed-v1 [217] Mistral Post Bi-Dir $\times$ LoRA Instruction WCL→SCL ZS Uni.
NV-Embed-v2 Mistral Post Bi-Dir $\times$ Unknown Instruction WCL→SCL ZS Uni.
Qwen Embedding (Alibaba)
GTE-Qwen2 [218] Qwen Special Last Bi-Dir $\times$ $\times$ Instruction WCL→SCL ZS Uni.
Qwen3 Embedding [219] Qwen Special Last Causal $\times$ $\times$ Instruction WCL→SCL→MG ZS Uni.
SFR-Embedding (Salesforce)
SFR-Mistral [220] Mistral Special Last Causal $\times$ LoRA Instruction SCL ZS Uni.
SFR-Embedding-2 Mistral Special Last Causal $\times$ $\times$ Instruction Unknown ZS Uni.
4.1 Backbone Selection
Model
The LLM-based embedders appearing in Table 4 use numerous open-source LLMs as backbones. Among the encoder-decoder backbones, T5 (GTR) has a dominant position, which is not surprising, as T5 and its variants have been leading innovations in encoder-decoder PLMs. Among decoder-only backbones, the most popular is Mistral (29 / 69), followed by LLaMA (21 / 69) and Qwen (15 / 69). The popularity of the Mistral is strongly connected to its excellent performance: a recent study [193] find that Mistral’s embedding performance will not drop significantly compared to the other LLMs when converting causal attention to bi-directional attention. This characteristic may derive from its pre-training or instruction fine-tuning phase, but no information has been revealed. Specifically, DIFFEMBED [202] employs the diffusion language model Dream 7B [221] as its backbone, demonstrating exceptional potential.
Parameter Size
4.2 Architecture Improvement
Denoting $d$ as the dimension of hidden states, $L$ as the number of transformer layers, and $\mathcal{V}$ as the vocabulary size, without loss of generality, we divide the LLM $F$ into multiple Transformer layers and a decoding layer, which is denoted as
$F=g\circ[f^{(L)}\circ\cdots\circ f^{(1)}]\circ f^{(0)}=g\circ f$
where $f^{(0)}:\mathbb{R}^{|\mathcal{V}|}\to\mathbb{R}^{d}$ is the input layer, $f^{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the $l$-th transformer layer and $g:\mathbb{R}^{d}\to\mathbb{R}^{|\mathcal{V}|}$ is the decoding layer. Although some studies [222, 173] have found that the optimal embeddings do not (solely) originate from the final layer, the vast majority of current approaches still derive embeddings from the pooling of hidden states from the final layer. We use $f$ to represent $f^{(L)}\circ\cdots\circ f^{(1)}]\circ f^{(0)}$ for convenience.
4.2.1 Traditional Pooling Strategy
Similar to traditional language models, the LLM includes a decoding layer that completes the mapping from the hidden state space to the token vocabulary space. To the best of our knowledge, almost all LLMs adhere to the original design of GPT [223], utilizing a simple, unbiased linear layer as the decoding layer. When LLMs are used as textual embedders, embeddings are obtained by a specific pooling strategy at hidden states, while mapping to the token space is no longer necessary. Therefore, most methods directly discard the decoding layer and opt for a pooling strategy on the hidden state of the last transformer layer’s output to obtain a single embedding for each text. To derive a universal formula, we introduce some optional components: (1) Instruction $\mathcal{I}$}; (2) In-Context Examples $\mathcal{E}$}; (3) Prompt Template prompt($\cdot$); (4) Special Token Sequences $\mathcal{S}$}. Then we can express the output of the LLM-based embedder as
$[{\bm{h}}^{(L)}_{0},{\bm{h}}^{(L)}_{{\mathcal{I}}_{1}},\cdots{\bm{h}}^{(L)}_{{\mathcal{I}}_{m}},{\bm{h}}^{(L)}_{{\mathcal{E}}_{1}},\cdots{\bm{h}}^{(L)}_{{\mathcal{E}}_{t}},{\bm{h}}^{(L)}_{x_{1}},\cdots{\bm{h}}^{(L)}_{x_{n}},{\bm{h}}^{(L)}_{{\mathcal{S}}_{1}},\cdots,{\bm{h}}^{(L)}_{{\mathcal{S}}_{k}}]=\mathcal{H}=f({\mathcal{I}}\oplus{\mathcal{E}}\oplus{\rm prompt}(x)\oplus{\mathcal{S}}),$ (19)
where ${\bm{h}}^{(L)}$ is the $L$-layer hidden state of $f$; $m$, $t$, $n$, and $k$ denote the number of tokens in $\mathcal{I}$}, $\mathcal{E}$}, ${\rm prompt}(x)$, and $\mathcal{S}$}, respectively. Specially, ${\bm{h}}^{(L)}_{0}$ represents the last hidden state corresponding to the starting token (e.g., [<pad>] for T5 or [<s>] for LLaMA) added by the models by default. Although some works [153, 193, 197, 67] splice instructions or in-context examples in front of the input text, obviously, its hidden states are not considered during the pooling stage. Therefore, we can expressed the pooling strategy as $\mathcal{P}(\cdot)$:
[…]
where $n$ is the token number of the input $x$, depending on the tokenizer of the LLMs. In this section, we have categorized the pooling strategy into six distinct types: (1) First Pooling; (2) Mean Pooling; (3) Last Pooling; (4) Post-Interaction Pooling; (5) Multi-Layer Pooling; (6) Generative Pooling. Each method will be introduced below separately.